# Document Valuation in LLM Summaries: A Cluster Shapley Approach

### v2

@@ 4.2 Revenue Attribution
An important application of our Shapley-based document valuation framework is enabling fair and economically grounded revenue sharing between platforms and document owners. In practice, platforms generate revenue either at the query level – for instance, through per-query advertising or API charges – or through subscription-based models that aggregate revenue over a large number of queries. Our framework naturally accommodates both settings and provides a principled method for revenue allocation based on each document’s marginal contribution to the LLM-generated output.
In real-world applications, platforms running LLM-based summarization services can decide a revenue share rate $\beta\in[0,1]$ for distribution among documents while retaining $1-\beta$ revenue. For example, ProRata.AI provides an attribution solution in AI search and shares revenue 50/50 with its sources (prorata2025). However, it uses a proprietary algorithmic approach to score and determine attribution (prorataUMG2025), the details of which are not publicly known.
@@
Query-Level Revenue. In settings where revenue is generated on a per-query basis—such as through API billing or advertising impressions—the allocation is straightforward. For a given query $q$ that yields revenue $r_{q}$, the payment to document $i$ is:
@@
In sum, our method offers a transparent and theoretically grounded mechanism for sharing platform revenue with document providers. Unlike flat licensing or arbitrary heuristics—such as attributing value based solely on user clicks—our framework accounts for each document’s marginal utility within the LLM summarization process. This avoids two key issues with click-based attribution: (1) many users do not click through to original documents, and (2) click rates are strongly influenced by the position or prominence of document presentation. In contrast, Shapley values capture each document’s actual contribution to the generated output. This enables fair compensation aligned with true informational value and supports both real-time (query-level) and batch (aggregate) attribution.
@@ 4.3 Approximating Shapley Value: Cluster Shapley Approach
Computational Challenge: While Shapley valuation is a theoretically appealing construct, evaluating Shapley values for source attribution presents a significant computational challenge even in moderate-sized settings. The computational cost associated with Shapley calculation exhibits exponential complexity – if we have $S_{q}$ relevant documents for query $q$, the number of both summarization and evaluations scales needed to calculate Shapley values scale as $2^{|S_{q}|}-1$, where $|S_{q}|$ is the number of relevant documents. This rapid growth in the number of calculations makes exact Shapley computation infeasible for large datasets, as the number of summarizations and evaluations quickly becomes overwhelming. Essentially, for each combination of documents, we need the LLM to generate a new summary and then perform an evaluation of that summary. While parallelization and batch processing can reduce latency, the overall computational burden remains substantial. As discussed in a later section $\S$6.5.2, even for a query with eight relevant documents, exact Shapley computation involves processing 255 subsets, leading to significant API costs and delays in LLM settings. These constraints suggest that exact Shapley methods in large-scale applications (e.g., document valuation for large platforms using LLMs) are infeasible and we therefore need efficient approximation algorithms.
@@
Researchers have proposed a number of algorithms designed to address the computational challenge associated with Shapley calculations. These approaches typically adopt a variety of sampling techniques to reduce the computational cost associated with Shapley calculation. One widely used method is Monte Carlo algorithm (mann1960values), which estimates Shapley values by randomly sampling permutations and computing marginal contributions across these samples. While this approach reduces computational cost compared to exact Shapley, it still requires a large number of samples to achieve reasonable accuracy. Truncated Monte Carlo (ghorbani2019data) improves efficiency by stopping the calculation early when additional samples provide diminishing returns below a threshold, significantly cutting down computational overhead. Another popular approach, Kernel SHAP (lundberg2017unified), employs a regression-based approximation to estimate Shapley values. However, none of these approaches leverage the textual content of documents when approximating Shapley values, treating them purely as independent units.
@@
Key Idea: Motivated by this limitation, we propose a novel Cluster Shapley algorithm that integrates semantic information from text embeddings to improve efficiency while preserving accuracy. Instead of treating documents as independent, our method utilizes LLM-generated embeddings to identify and group similar documents, reducing redundant evaluations. Our core idea is intuitive: documents with similar content should have comparable contributions to the final summary and, therefore, should receive similar Shapley values. The key strength of our approach is that it leverages the textual content of the documents and the LLM’s numerical representation of this textual content (i.e., text embedding) to help approximate and simplify Shapley calculations.
@@
Text embedding techniques convert large chunks of text—such as sentences, paragraphs, or documents—into numerical vectors that capture semantic information. Earlier embedding methods, such as Word2Vec (mikolov2013efficient) and GloVe (pennington2014glove), are based on shallow neural networks and co-occurrence statistics, learning word-level embeddings by predicting surrounding context words or factorizing word co-occurrence matrices. These embeddings typically represent each word with a fixed vector, independent of context. In contrast, modern LLM-based embeddings, such as those produced by OpenAI’s latest text-embedding models, are generated using Transformer-based architectures and are pretrained on massive text corpora via next-token prediction objectives. These newer embeddings are contextualized—meaning the vector for a word or sentence depends on its surrounding context—and are typically high-dimensional (e.g., 3072 dimensions in OpenAI’s text-embedding-3-large model). Unlike generative LLMs designed for tasks like chat or text generation, embedding models are optimized to produce semantically meaningful representations suitable for a wide range of downstream tasks. LLM-based embedding vectors have been successfully applied to a wide variety of discriminative tasks, including text classification, document retrieval, sentiment analysis, and predicting the attractiveness of news headlines (patil2023survey; ye2025lola). In our setting, we leverage these embeddings to cluster similar documents before computing Shapley values, allowing us to reduce redundant calculations.
@@
Our Approach: We outline our proposed Cluster Shapley Algorithm in Algorithm 1. Cluster Shapley begins with a preprocessing step, where for a given query $q$, we first determine the set of relevant documents $S_{q}\subseteq D$. This retrieval step ensures that only contextually relevant documents are considered; e.g., if the query pertains to political news, unrelated sports articles will be excluded from summarization. In $\S$6.2, we discuss the retrieval and summarization steps, and related literature in further detail. For each document $i\in S_{q}$, we also obtain its embedding vector $e_{i}$, using an LLM embedding model. This step can be performed using proprietary models like OpenAI and Gemini or open-source alternatives such as Llama/Alpaca.
@@
4.3.1 Step 1: Document Clustering with the Distance Constraint
[…]
Algorithm 2 Adaptive Distance-Constrained DBSCAN
  Run the standard DBSCAN with $\texttt{MinPts}=1$:
  • For each document $i$: find $r$-neighborhood $N_{r}(i)=\{j:M_{ij}\leq r\}$.
  • Connect points that are within $r$ distance.
  Check the distance constraint:
end while
Lastly, we discuss the choice of the clustering diameter $\epsilon$, a key hyperparameter that governs both the approximation error and the computational cost. Smaller values of $\epsilon$ lead to more accurate approximations of the Shapley value but incur higher computational cost, as formalized in §4.4. In practice, one can perform standard hyperparameter tuning: use a separate dataset – such as one from historical logs – to select the value of $\epsilon$ that best balances approximation accuracy and computational efficiency. This selected value can then be fixed for future use. Later in $\S$7.2, we empirically show how $\epsilon$ governs the trade-offs between computational costs and approximation error in our application setting.
4.3.2 Steps 2 and 3: Cluster-level Shapley Value Computation and Document Value Allocation
@@
Note that our Cluster Shapley framework is flexible with respect to the choice of Shapley computation method. In Step 2, any Shapley value algorithm can be applied, including exact calculations or approximation techniques. In case we have a very large dataset or if there are a large number of clusters, approximation methods such as Monte Carlo and Kernel SHAP can be employed to further reduce computational costs while maintaining reasonable accuracy. We formally discuss this additional approximation in $\S$4.4.3.
@@
In summary, our Cluster Shapley Algorithm offers two advantages. First, it reduces complexity by grouping similar documents into clusters, which decreases the number of subsets to process from $2^{|S_{q}|}-1$ to $2^{m}-1$ ($m$ is the number of clusters), thereby significantly reducing computation time for summarization and evaluation. Second, it provides flexibility in balancing speed and precision through the adjustable $\epsilon$ parameter. For tasks requiring high precision, a smaller $\epsilon$ can be selected, while tasks prioritizing speed can utilize a larger $\epsilon$.
@@ 4.4 Theoretical Analysis of Cluster Shapley Algorithm
We now develop theoretical guarantees for our proposed Cluster Shapley algorithm. First, in $\S$4.4.1, we show that the approximation error of Cluster Shapley can be bounded. Second, we characterize the computational complexity of the algorithm in $\S$4.4.2. Finally, in $\S$4.4.3, we consider a setting where $m$ is still large (so that $2^{m}$ complexity is burdensome). In this case, we show that it is feasible to use an approximation algorithm (e.g., Monte Carlo) in the second step of Cluster Shapley and theoretically bound the approximation error.
@@ 4.4.1 Approximation Error of Cluster Shapley
If a cluster $G_{k}$ is such that all member documents are nearly identical, i.e., the maximum difference $\phi_{\max}-\phi_{\min}=\delta\ll 1$), then each document’s Cluster Shapley value is within $\delta$ of its true Shapley value, $|\hat{\phi}_{i}-\phi_{i}|\leq\delta$ for all $i\in G_{k}$. In particular, if documents in $G_{k}$ are symmetric (meaning $v(S\cup\{i\})=v(S\cup\{j\})$ for all $S$ and all $i,j\in G_{k}$), then $\hat{\phi}_{i}=\phi_{i}$ exactly for all $i\in G_{k}$.
@@ 4.4.2 Computational Complexity Analysis
    Document clustering complexity: The algorithm must compute embeddings for $n$ documents and then cluster them based on pairwise distances. Generating embeddings (using a pre-trained LLM embedding model) takes $O(n)$ operations. Computing the pairwise distance matrix naively costs $O(n^{2})$ time (since there are $\binom{n}{2}$ pairs). The clustering itself using DBSCAN will typically require examining each point’s neighbors; DBSCAN in worst-case can be $O(n^{2})$, though it can approach $O(n\log n)$ on average (ester1996density). In summary, the clustering step is polynomial in $n$. We can reasonably approximate it as $O(n^{2})$ in the worst case, which is manageable for moderate $n$. This phase is far more efficient than an exhaustive Shapley computation (which is exponential in $n$).
@@
Combining the two phases, the overall worst-case time complexity is $O(n^{2})+O(2^{m})$. It is important to note that the exponential complexity in $m$ is unavoidable if using the exact calculation in Step 2 of the Cluster Shapley algorithm. Cluster Shapley does not magically circumvent the combinatorial explosion; rather, it reduces the problem size from $n$ to $m$ by leveraging redundancy in the document. In scenarios where $n$ is large but the effective number of independent information sources is small (many documents are repeats or very similar), this yields a drastic speed-up. In the best case, if all $n$ documents were near-duplicates of a few types, $m$ would be small (say $m=O(\log n)$ or even $O(1)$), and then the overall algorithm runs in polynomial time $O(n^{2}+2^{O(\log n)})=O(n^{2}+n^{c})$ (or even $O(n^{2})$), which is efficient. In the worst case, where no two documents are alike ($m=n$), the complexity $O(n^{2}+2^{n})$ reverts to that of exact Shapley.
@@ 4.4.3 Large $m$ Scenario
In some cases, the number of clusters ($m$) can be large. Then, computing and evaluating all $2^{m}$ subsets is burdensome. In such cases, Cluster Shapley can be combined with any cluster-level Shapley approximation algorithm. Specifically, instead of exactly evaluating $v_{\mathcal{G}}$ on all cluster subsets, we can apply one of the existing Shapley approximation algorithms $\mathcal{A}$ with known convergence rate, approximation error, and computational cost. We formally describe this generalized procedure in Algorithm 3, and establish the corresponding error bound and computational complexity in Theorem 2.
@@
Suppose (i) Assumption 1 holds, (i) Approximation algorithm $\mathcal{A}$ ensures that each cluster Shapley value $\tilde{\phi}_{G_{k}}$ satisfies $|\tilde{\phi}_{G_{k}}-\hat{\phi}_{G_{k}}|\leq\epsilon_{\mathcal{A}}(\eta)$ with probability at least $1-\eta$, and (iii) Approximation algorithm $\mathcal{A}$ has computational complexity $C_{\mathcal{A}}(m)$, then for any document $i\in G_{k}$, with probability at least $1-\eta$:
@@
Suppose $\mathcal{A}$ is Monte Carlo approximation with $N$ random permutations. Under Assumption 1 and bounded marginal contributions $V_{\max}$, for any document $i\in G_{k}$ and probability $1-\eta$, we have:
@@
In sum, our theoretical analysis shows that Cluster Shapley is a principled, flexible algorithm that offers provable accuracy-efficiency trade-offs. We now demonstrate the empirical performance of our algorithm in the following sections.
@@ 5 Application Setting: Amazon Review Dataset
We now present an application of our algorithm to a real setting. We use the publicly available Amazon Product Reviews dataset as the empirical context to demonstrate the performance of our document valuation approach. This dataset was collected by hou2024bridging and has been extensively utilized in recent research studies on a variety of topics, including sentiment analysis (haque2018sentiment), sequential product search and recommendation (hou2024bridging), fine-tuning of LLMs (zhang2024llasa), and evaluation of LLM alignment (shankar2024validates). The dataset spans from May 1996 to September 2023, featuring over 571.54 million reviews from 54.51 million users and covering 48.19 million unique items. It is organized into 33 distinct categories, including electronics, household goods, clothing, and books. This user review data set consists of textual feedback provided by users that captures their opinions, ratings, and experiences with products. A comprehensive analysis of review categories, basic statistics, and detailed data field information is available in hou2024bridging.
While it is well-established that review valence and content can help consumers make better decisions (chevalier_mayzlin_2006), it is also well-understood that it is hard for consumers to process the large amounts of information/text in reviews. For example, the most popular products in the data have hundreds or even thousands of reviews. Thus, consumers must often sift through hundreds of reviews to extract relevant insights. This information overload makes it difficult for users to efficiently locate specific details (e.g., product quality, value for money, durability, ease of return).
@@
To compute the Shapley value of each individual review (within the context of a given product), we need to first specify the distribution of queries, $g(q)$, that consumers use when requesting summaries for this product. While this query distribution is not publicly available, for each product, we observe the set of popular attributes frequently mentioned by customers. For instance, in Figure 1, for the wireless controller product, attributes such as “Functionality” and “Controller quality” are among the most common aspects of the product that users are concerned about. This information allows us to craft a proxy distribution of user queries for each product that mimics the real distribution of queries. Specifically, for each product, we design two queries based on the top two attributes most frequently mentioned in its reviews, as shown in the last column of Table LABEL:tab:designed_queries_full. Each query is assigned an equal probability, contributing equally to the overall valuation. [^7] In practice, the platform will have data on the true distribution of queries and can directly use that empirical distribution in their analyses. As discussed earlier, our approach is agnostic to the exact distribution of queries.

### v3

@@ 4.2 Revenue Attribution
An important application of our Shapley-based document valuation framework is enabling fair and economically grounded revenue sharing between platforms and document owners. In practice, platforms generate revenue either at the query level, for instance through per-query advertising or API charges, or through subscription-based models that aggregate revenue over a large number of queries. Our framework can accommodate both settings and provide a principled method for revenue allocation based on each document’s marginal contribution to the LLM-generated output.
To illustrate how such a framework can be operationalized in practice, consider platforms running LLM-based summarization services. Such platforms can determine a revenue share rate $\beta\in[0,1]$ for distribution among documents while retaining $1-\beta$ as platform revenue. For example, ProRata.AI provides an attribution solution in AI search and shares revenue 50/50 with its sources (51). However, it uses a proprietary algorithmic approach to score and determine attribution (66), the details of which are not publicly known.
@@
Query-Level Revenue. In settings where revenue is generated on a per-query basis-such as through API billing or advertising impressions-the allocation is straightforward. For a given query $q$ that yields revenue $r_{q}$, the payment to document $i$ is:
@@
In sum, our method offers a transparent and theoretically grounded mechanism for sharing platform revenue with document providers. Unlike flat licensing or arbitrary heuristics-such as attributing value based solely on user clicks-our framework accounts for each document’s marginal utility within the LLM summarization process. This avoids two key issues with click-based attribution: (1) many users do not click through to original documents, and (2) click rates are strongly influenced by the position or prominence of document presentation. In contrast, Shapley values capture each document’s actual contribution to the generated output. This enables fair compensation aligned with true informational value and supports both real-time (query-level) and batch (aggregate) attribution.
@@ 4.3 Approximating Shapley Value: Cluster Shapley Approach
Computational Challenge: While Shapley valuation is a theoretically appealing construct, evaluating Shapley values for source attribution presents a significant computational challenge even in moderate-sized settings. The computational cost associated with Shapley calculation exhibits exponential complexity – if we have $S_{q}$ relevant documents for query $q$, the number of summarizations and evaluations needed to compute Shapley values scales as $2^{|S_{q}|}-1$, where $|S_{q}|$ is the number of relevant documents. This rapid growth in the number of calculations makes exact Shapley computation infeasible for large datasets, as the number of summarizations and evaluations quickly becomes overwhelming. Essentially, for each combination of documents, we need the LLM to generate a new summary and then perform an evaluation of that summary. These constraints suggest that exact Shapley methods in large-scale applications (e.g., document valuation for large platforms using LLMs) are infeasible, and we therefore need efficient approximation algorithms.
@@
Researchers have proposed a number of algorithms designed to address the computational challenge associated with Shapley calculations. These approaches typically adopt a variety of sampling techniques to reduce the computational cost associated with Shapley calculation. One widely used method is Monte Carlo algorithm (42), which estimates Shapley values by randomly sampling permutations and computing marginal contributions across these samples. While this approach reduces computational costs compared to exact Shapley values, it still requires a large number of samples to achieve reasonable accuracy. Truncated Monte Carlo (20) improves efficiency by stopping the calculation early when additional samples provide diminishing returns below a threshold, significantly cutting down computational overhead. Another popular approach, Kernel SHAP (40), employs a regression-based approximation to estimate Shapley values. However, none of these approaches leverage the textual content of documents when approximating Shapley values, treating them purely as independent units.
@@
Key Idea: Motivated by this limitation, we propose the Cluster Shapley algorithm. Instead of treating documents as independent, our method utilizes LLM-generated embeddings to identify and group similar documents, reducing redundant evaluations. Our core idea is intuitive: documents with similar content should have comparable contributions to the final summary and, therefore, should receive similar Shapley values. Our approach leverages the textual content of the documents and the LLM’s numerical representation of this textual content (i.e., text embedding) to help approximate and simplify Shapley calculations.
@@
Text embedding techniques convert large chunks of text-such as sentences, paragraphs, or documents-into numerical vectors that capture semantic information. Earlier embedding methods, such as Word2Vec (46) and GloVe (49), are based on shallow neural networks and co-occurrence statistics, learning word-level embeddings by predicting surrounding context words or factorizing word co-occurrence matrices. These embeddings typically represent each word with a fixed vector, independent of context. In contrast, modern LLM-based embeddings, such as those produced by OpenAI’s latest text-embedding models, are generated using Transformer-based architectures and are pretrained on massive text corpora via next-token prediction objectives. These newer embeddings are contextualized-meaning the vector for a word or sentence depends on its surrounding context-and are typically high-dimensional (e.g., 3072 dimensions in OpenAI’s text-embedding-3-large model). Unlike generative LLMs designed for tasks like chat or text generation, embedding models are optimized to produce semantically meaningful representations suitable for a wide range of downstream tasks. LLM-based embedding vectors have been successfully applied to a wide variety of discriminative tasks, including text classification, document retrieval, sentiment analysis, and predicting the attractiveness of news headlines (48; 69). In our setting, we leverage these embeddings to cluster similar documents before computing Shapley values, allowing us to reduce redundant calculations.
@@
Our Approach: We outline our proposed Cluster Shapley Algorithm in Algorithm 1. Cluster Shapley begins with a preprocessing step, where for a given query $q$, we first determine the set of relevant documents $S_{q}\subseteq D$. This retrieval step ensures that only contextually relevant documents are considered; e.g., if the query pertains to political news, unrelated sports articles will be excluded from summarization. In $\S$6.1, we discuss the retrieval and summarization steps, and related literature in further detail. For each document $i\in S_{q}$, we also obtain its embedding vector $e_{i}$, using an LLM embedding model. This step can be performed using proprietary models like OpenAI and Gemini or open-source alternatives such as Llama/Alpaca.
@@
The goal of Step 1 is to partition the documents into $m$ non-overlapping clusters, i.e., $D=\bigcup_{k=1}^{m}G_{i}$ with $G_{i}\cap G_{j}=\emptyset$ for $i\neq j$. We can employ any off-the-shelf clustering algorithms in Step 1, such as K-Means (39) or Density-Based Spatial Clustering of Applications with Noise (DBSCAN) (15). However, standard clustering algorithms do not always enforce uniform intra-cluster proximity-that is, they may allow pairs of documents within the same cluster to be far apart, particularly in the presence of high noise or uneven density. To address this limitation, we propose an adaptive DBSCAN algorithm (Algorithm 3) detailed in Web Appendix §B.1. For comparison, we also empirically evaluate the standard DBSCAN in §7.3, which performs worse in our setting.
The clustering diameter $\epsilon$ is a key hyperparameter that governs both the approximation error and the computational cost. Smaller values of $\epsilon$ lead to more accurate approximations of the Shapley value but incur higher computational cost, as formalized in §4.4. In practice, one can perform standard hyperparameter tuning: use a separate dataset to select the value of $\epsilon$ that best balances approximation accuracy and computational efficiency. This selected value can then be fixed for future use. Later in $\S$7.1, we empirically show how $\epsilon$ governs the trade-offs between computational costs and approximation error in our application setting.
@@
Note that our Cluster Shapley framework is flexible with respect to the choice of Shapley computation method. In Step 2, any Shapley value algorithm can be applied, including exact calculations or approximation techniques. In case we have a very large dataset or if there are a large number of clusters, approximation methods such as Monte Carlo can be employed to further reduce computational costs while maintaining reasonable accuracy. We formally discuss this additional approximation in $\S$4.4.3.
@@ 4.4 Theoretical Analysis of Cluster Shapley Algorithm
We now analyze our proposed Cluster Shapley algorithm. First, in $\S$4.4.1, we show that the approximation error of Cluster Shapley can be bounded. Second, we characterize the computational complexity of the algorithm in $\S$4.4.2. Finally, in $\S$4.4.3, we consider a setting where $m$ is still large (so that $2^{m}$ complexity is burdensome). In this case, we show that it is feasible to use an approximation algorithm (e.g., Monte Carlo) in the second step of Cluster Shapley and bound the approximation error.
@@ 4.4.1 Approximation Error of Cluster Shapley
If a cluster $G_{k}$ is such that all member documents are nearly identical, i.e., the maximum difference $\phi_{\max}-\phi_{\min}=\delta\ll 1$, then each document’s Cluster Shapley value is within $\delta$ of its true Shapley value, $|\hat{\phi}_{i}-\phi_{i}|\leq\delta$ for all $i\in G_{k}$. In particular, if documents in $G_{k}$ are symmetric (meaning $v(S\cup\{i\})=v(S\cup\{j\})$ for all $S$ and all $i,j\in G_{k}$), then $\hat{\phi}_{i}=\phi_{i}$ exactly for all $i\in G_{k}$.
@@ 4.4.2 Computational Complexity Analysis
    Document clustering complexity: The algorithm must compute embeddings for $n$ documents and then cluster them based on pairwise distances. Generating embeddings (using a pre-trained LLM embedding model) takes $O(n)$ operations. Computing the pairwise distance matrix naively costs $O(n^{2})$ time (since there are $\binom{n}{2}$ pairs). The clustering itself using DBSCAN will typically require examining each point’s neighbors; DBSCAN in worst-case can be $O(n^{2})$, though it can approach $O(n\log n)$ on average (15). In summary, the clustering step is polynomial in $n$. We can reasonably approximate it as $O(n^{2})$ in the worst case, which is manageable for moderate $n$. This phase is far more efficient than an exhaustive Shapley computation (which is exponential in $n$).
@@
Combining the two phases, the overall worst-case time complexity is $O(n^{2})+O(2^{m})$. It is important to note that the exponential complexity in $m$ is unavoidable if using the exact calculation in Step 2 of the Cluster Shapley algorithm. Cluster Shapley does not magically circumvent the combinatorial explosion; rather, it reduces the problem size from $n$ to $m$ by leveraging redundancy in the document.
@@ 4.4.3 Large $m$ Scenario
In some cases, the number of clusters, $m$, can be large. Then, computing and evaluating all $2^{m}$ subsets is burdensome. In such cases, Cluster Shapley can be combined with any cluster-level Shapley approximation algorithm. Specifically, instead of exactly evaluating $v_{\mathcal{G}}$ on all cluster subsets, we can apply one of the existing Shapley approximation algorithms $\mathcal{A}$ with known convergence rate, approximation error, and computational cost. We formally describe this generalized procedure in Algorithm 2, and establish the corresponding error bound and computational complexity in Theorem 2.
@@
Suppose (i) Assumption 1 holds, (ii) Approximation algorithm $\mathcal{A}$ ensures that each cluster Shapley value $\tilde{\phi}_{G_{k}}$ satisfies $|\tilde{\phi}_{G_{k}}-\hat{\phi}_{G_{k}}|\leq\epsilon_{\mathcal{A}}(\eta)$ with probability at least $1-\eta$, and (iii) Approximation algorithm $\mathcal{A}$ has computational complexity $C_{\mathcal{A}}(m)$, then for any document $i\in G_{k}$, with probability at least $1-\eta$:
@@
Suppose $\mathcal{A}$ is the Monte Carlo approximation algorithm with $N$ random permutations. Under Assumption 1 and bounded marginal contributions $V_{\max}$, for any document $i\in G_{k}$ with probability $1-\eta$, we have:
@@
In sum, our theoretical analysis shows that Cluster Shapley is a principled, flexible algorithm that offers provable efficiency-accuracy trade-offs. We now demonstrate the empirical performance of our algorithm in the following sections.
@@ 5 Application Setting: Amazon Review Dataset
We now present an application of our algorithm to a real setting. We use the publicly available Amazon Product Reviews dataset as the empirical context to demonstrate the performance of our document valuation approach. This dataset was collected by Hou et al. (29) and has been used extensively in recent research studies on a variety of topics, including sentiment analysis (28), sequential product search and recommendation (29), fine-tuning of LLMs (73), and evaluation of LLM alignment (61). The dataset spans from May 1996 to September 2023, featuring over 571.54 million reviews from 54.51 million users and covering 48.19 million unique items. It is organized into 33 distinct categories, including electronics, household goods, clothing, and books. This user review data set consists of textual feedback provided by users that captures their opinions, ratings, and experiences with products. A comprehensive analysis of review categories, basic statistics, and detailed data field information is available in Hou et al. (29).
While it is well-established that review valence and content can help consumers make better decisions (9), it is also well-understood that it is hard for consumers to process the large amounts of information/text in reviews. For example, the most popular products in the data have hundreds or even thousands of reviews. Thus, consumers must often sift through hundreds of reviews to extract relevant insights. This information overload makes it difficult for users to efficiently locate specific details (e.g., product quality, value for money, durability, ease of return).
@@
To compute the Shapley value of each individual review (within the context of a given product), we need to first specify the distribution of queries, $g(q)$, that consumers use when requesting summaries for this product. While this query distribution is not publicly available, for each product, we observe the set of popular attributes frequently mentioned by customers. For instance, in Figure 1, for the wireless controller product, attributes such as “Functionality” and “Controller quality” are among the most common aspects of the product that users are concerned about. This information allows us to craft a proxy distribution of user queries for each product that mimics the real distribution of queries. Specifically, for each product, we design two queries based on the top two attributes most frequently mentioned in its reviews, as shown in the last column of Table LABEL:tab:designed_queries_full. Each query is assigned an equal probability, contributing equally to the overall valuation. [^8] In practice, the platform will have data on the true distribution of queries and can directly use that empirical distribution in its analyses. As discussed earlier, our approach is agnostic to the exact distribution of queries.